OpenStudy (anonymous):

If f(x) = x + 4 and g(x) = 2x - 3, find (g - f)(2).

OpenStudy (anonymous):

\( (g-f)(2) = g(2) - f(2)\) Can you do the next step?

OpenStudy (anonymous):

No .. im sorry i really dont know this . @OneMathCat

OpenStudy (anonymous):

\(g(x) = 2x - 3\) \(g(2) = 2\cdot 2 - 3\) Do you see the pattern?

hero (hero):

@ onemathcat, you have to be careful with the way you explain things. Let's assume the user understands everything you explain to her , but then she goes to school the next day, and even though She has the correct Answer and all the steps, the teacher proceeds to ask her, What is (g-f)(x)? With your limited explanation, she may not be able to answer the teacher's question.

OpenStudy (anonymous):

this is an online class , she doesnt ask anything . Thank you for saying that though @Hero . I dont understand though ..

OpenStudy (anonymous):

@Hero Hi ... I certainly wasn't intending to stop there! But, there was no one in sight ... evidently the user had gone offline. What is the recommendation at that point? Should you then type out the complete answer? I know the code of honor is certainly NOT just to give the answer...

hero (hero):

Not every student here is taking online classes.

OpenStudy (anonymous):

but i am , and i still dont understand ..

hero (hero):

What gm trying to suggest here is to Compute (g-f)(x) first before computing (g-f)(2)

OpenStudy (anonymous):

When you see \(g(x) = 2x - 3\), it's what is called "function notation". Think of it this way: g(blah) = 2(blah) - 3. So: g(1) = 2(1) - 3 g(2) = 2(2) - 3 g(10.7) = 2(10.7) - 3 Are you seeing the pattern?

OpenStudy (anonymous):

where did 1 come from ?

OpenStudy (anonymous):

and 2?

OpenStudy (anonymous):

@Hero I'm happy to have you take over. No problem. Have a great day!

hero (hero):

I don't want to takeover.

OpenStudy (anonymous):

did someone get 2x+2 ??

OpenStudy (anonymous):

@Hero

hero (hero):

Explain how you got 2x+2

OpenStudy (anonymous):

2x-3 - 4x+4 which gave me x+1 and i multiplied that by 2

hero (hero):

Somehow you've managed to confuse yourself

OpenStudy (anonymous):

but thats what it said right